Amina sells T-shirts and singlets. A T-shirt costs 5times as much as a singlet. For GH¢800.00, she can buy 32 more singlets than a T-shirt. How much does a T-shirt cost?

Answer by jorel1380(3719)

Let n be cost of a singlet, and 5n be the cost of a T-shirt. Then:
800/5n=800/n -32
800=4000-160n
160n=3200
n=20
The cost of a singlet is GH 20; A T-shirt costs $100
☺☺☺☺